heathergreen:

What is the solution to the equation 5 (x minus 6) = 2 (x + 3)? x =

Shadow:

\[5(x - 6) = 2(x + 3)\] Distribute on both sides \[5x - 30 = 2x + 6\] Add 30 to both sides \[5x = 2x + 36\] Subtract 2x from both sides \[3x = 36\] Divide both sides by 3 \[x = 12\]

heathergreen:

thanks